In what distance will a car skid to a stop on a dry concrete road (µk = 0.7) if its brakes are locked when it is moving at 120 kph?


1
Expert's answer
2021-01-29T13:22:39-0500

Answer

Distance can be written as

"s=\frac{v^2}{2\mu g}=\frac{(33.3)^2}{2\times0.7\times9.8}=80.8m"